OpenStudy (anonymous):

(Must show work!) 1. During a car trip, the Webers buy gasoline at three different stations. At the first station, they pay $18.64 for 9 gallons of gas. at the second, they pay $28.54 for 14 gallons. At the third, they pay $33.44 for 16 gallons. Which station offers the lowest price per gallon?

OpenStudy (anonymous):

to find price (or $) per gallon, you simply divide the price ($) by the number of gallons. Station 1 price/gal =$18.64/9; Station 2 =28.54/14; and Station 3 = 33.54/16. Which one is the lowest?

OpenStudy (anonymous):

@michaelsmarketplace Is it the third one, right?